Consider again the problem of a car traveling along a banked turn. Sometimes roads have a "reversed" banking angle. That is, the road is tilted "away" from the center of curvature of the road. If the coefficient of static friction between the tires and the road is μs = 0.4, the radius of curvature is 25 m, and the banking angle is 14°, what is the maximum speed at which a car can safely navigate such a turn?
Answer #1

Solution) R = 25 m

Us = 0.4

Angle , theeta = 14°

V = ?

For safe turn

fs = (m(V^2))/(R)

fs = (Us)mg

(Us)mg = (m(V^2))/(R)

V^2 = (Us)(R)(g)

V^2 = (0.4)(25)(9.8)

V^2 = 98

V = (98)^(1/2)

V = 9.89 m/s